OpenStudy (anonymous):

integrate (x^3+x^4tanx) dx from x=-pi/4 to pi/4

OpenStudy (anonymous):

\[\int\limits_{-\pi/4}^{\pi/4} (x^3+x^4tanx) dx\]

hartnn (hartnn):

check whether x^3+x^4 tan x is odd function or not. if its an odd function, then the integral =0

hartnn (hartnn):

the function is odd if f(x) = -f(-x) and a property of indefinite integral is \(\int \limits_{-a}^af(x)dx=0\) if f(x) is odd function, \(\int \limits_{-a}^af(x)dx=2\int \limits_{0}^af(x)dx\) if f(x) is even function.

OpenStudy (anonymous):

what if its neither lol

hartnn (hartnn):

if its neither, you have to integrate the function as it is....no simplification...